Solve for x (round to the nearest thousandth): 8+(2x)^3=1000
Answer by lwsshak3(11628)

Solve for x (round to the nearest thousandth): 8+(2x)^3=1000
**
8+(2x)^3=1000
(2x)^3=1000-8
8x^3=992
x^3=992/8=124
use calculator to take cube root of both sides
x≈4.987